OpenStudy (jmoores):

Snowy's Snow Cones has a special bubble gum snow cone on sale. The cone is a regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. The radius of the snow cone is 3 inches, and the height of the cone is 5 inches. If the diameter of the bubble gum is 1.1 inches, which of the following can be used to calculate the volume of the cone that can be filled with flavored ice? (5 points) Select one: a. 1 over 3 (3.14)(32)(5) − 4 over 3 (3.14)(1.13) b. 1 over 3 (3.14)(52)(3) − 4 over 3 (3.14)(1.13) c. 1 over 3 (3.14)(32)(5) − 4 over 3 (3.14)(0.553) d. 1 over 3

OpenStudy (aakashtomar):

Sorry for the delay. I was on a call, really very sorry.

OpenStudy (jmoores):

Don't worry, you're completely fine! :)

OpenStudy (jmoores):

I'm just so grateful I can get some kind of help honestly!

OpenStudy (aakashtomar):

If you subtract the volume of the sphere from the cone, you'll get the volume you can fill with flavored ice. Right?

OpenStudy (jmoores):

I think so

OpenStudy (aakashtomar):

Okay, let's make this better through a drawing.  My drawing is really bad, sorry, but I think you can visualize the chewing gum inside the cone.

OpenStudy (jmoores):

Ok, I see

OpenStudy (aakashtomar):

I mean, that is as far as I can infer from the question. Now, \[Volume (sphere) = 4/3 \pi r^3 \] \[Volume(cone) = 1/3 \pi r^2 \times h\]

OpenStudy (aakashtomar):

So that means, The volume left to put in flavored ice is : \[1/3 \pi r^2 - 4/3 \pi r^3\]

OpenStudy (aakashtomar):

Sorry, that is : \[1/3 \pi r^2 h - 4/3 \pi r^3\]

OpenStudy (aakashtomar):

So putting in the value for radius and height of the cone and radius of the sphere in the 2nd term will do the thing, I think.

OpenStudy (jmoores):

Ok, hold one, let me do it and then tell you what I get

OpenStudy (jmoores):

on*

OpenStudy (aakashtomar):

Sure!

OpenStudy (jmoores):

So wait, would I plug it in like this: 1/3(pi)3^2(5)-4/3)(pi)0.55)^2?

OpenStudy (aakashtomar):

Just that the radius of sphere is cubed, not squared.

OpenStudy (jmoores):

Ok, so switch it to cubed and then I should get my answer?

OpenStudy (aakashtomar):

I'm trying to see just that. I seriously think that this is the correct way to solve, but the problem is, the square of the radius of the cone isn't matching any options. It should be 9, but it's given 32 and 52 in the options. Also they haven't cubed the radius of the sphere, which is really wrong.

OpenStudy (jmoores):

The course I'm taking is awful, I can say that. Can you type the formula you think I need to use with the numbers plugged in so I have an idea of where you are getting at?

OpenStudy (jmoores):

I'm just confusing myself over here

OpenStudy (aakashtomar):

\[1/3 \times 3.14 \times (3)^2 \times 5 - 4/3 \times 3.14 \times (0.55)^3\]

OpenStudy (jmoores):

Oh, I see. Hold on, let me calculate it and tell you what I get.

OpenStudy (aakashtomar):

Oh! I See!

OpenStudy (aakashtomar):

Okay, I think they didn't raise the power but wrote the exponent beside the number. It must be c. They wrote 32 so as to say 3^2 and wrote 0.553 to say (0.55)^3

OpenStudy (jmoores):

Oh my gosh! I think you're right!

OpenStudy (aakashtomar):

Yeah, so the answer is c.

OpenStudy (jmoores):

That makes more sense now! Thank you so much! You were so so helpful! :)

OpenStudy (aakashtomar):

Nah, I'm always so happy to help. :) Welcome!